# A Rapid Trajectory Optimization and Control Framework for Resource-Constrained Applications

### v1

@@ II-D State and Input Constraints
With these modifications, the quadratic program of Eq. (31) is written with equality and inequality constraints as
$\min_{\bm{\chi}}J$ $=\frac{1}{2}\bm{\chi}^{\top}H\bm{\chi}+\bm{f}^{\top}\bm{\chi}$ (34)
[…]
$A\bm{\chi}$ $\leq\mathbf{b}$ (36)
where
$\bm{\chi}$ $=\begin{bmatrix}\bm{\alpha}\\ \varepsilon\end{bmatrix}\in\mathbb{R}^{(n+2)\times 1}$ (37)
$H$ $=\begin{bmatrix}H&\bm{0}_{n+1,1}\\ \bm{0}_{1,n+1}&\rho\end{bmatrix}\in\mathbb{R}^{(n+2)\times(n+2)}$ (38)
$A_{\text{eq}}$ $=\begin{bmatrix}\bm{\gamma}(-1)&0\\ \bm{\beta}(-1)&0\end{bmatrix}\in\mathbb{R}^{2\times(n+2)}$ (39)
$\mathbf{b}_{\text{eq}}$ $=\begin{bmatrix}0\\ 0\end{bmatrix}\in\mathbb{R}^{2\times 1}$ (40)
$A$ $=\begin{bmatrix}\mathcal{T}(\bm{\tau})&-V_{u}\\ -\mathcal{T}(\bm{\tau})&-V_{u}\\ -\bm{\beta}(\bm{\tau})&-V_{x^{\prime}}\\ -\bm{\beta}(\bm{\tau})&-V_{x^{\prime}}\end{bmatrix}\in\mathbb{R}^{4(n+1)\times(n+2)}$ (41)
$\mathbf{b}$ $=\begin{bmatrix}u_{\text{max}}^{\prime}\\ -u_{\text{min}}^{\prime}\\ x_{\text{max}}^{\prime}-x^{\prime}(-1)\\ -x_{\text{min}}^{\prime}+x^{\prime}(-1)\end{bmatrix}\in\mathbb{R}^{4(n+1)\times 1}$ (42)
and $\bm{f}$ remains unchanged from Eq. (33). Here, $\rho$ is the weight factor associated with the slack variable $\varepsilon$, $u_{\text{max}}^{\prime}/u_{\text{min}}^{\prime}$ and $x_{\text{max}}^{\prime}/x_{\text{min}}^{\prime}$ are the constraints on the control and velocity converted to the Chebyshev domain, and $V_{u}$ and $V_{x^{\prime}}$ are used to regulate the softness of the constraints. A comprehensive comparison of MPC3 with discrete MPC along with dimensions of the optimization program is given in Table I.
@@ III Performance Comparison
The constrained optimization problem, now formulated in the computational domain as a quadratic program, is solved with MATLAB®’s quadprog using the active-set algorithm and warm-start enabled. Importantly, note that of the matrices and vectors defined in the previous section, $H$, $A_{\text{eq}}$, $\mathbf{b}_{\text{eq}}$, and $A$ remain unchanged, regardless of time horizon or initial conditions. Only $\bm{f}$ and $\mathbf{b}$ need to be updated at each step, resulting in an efficient formulation. The coefficients and slack variable are found by minimizing the objective function of Eq. (34), and the instantaneous control is calculated as $u(-1)=\bm{T}(-1)\bm{\alpha}$. ICC, or any other integrator, can then be used to propagate the true dynamics until reaching the next control input time.
@@ III-A Variation with State Dimension
To establish the computational gains of the proposed formulation, extensive simulation studies have been conducted and their results are presented in Fig. 3. For the first analysis, the control horizon has been kept fixed at $p=5$ samples and the dimension of the state $q$ has been varied. The run time and memory requirement of a single iteration of the optimization is recorded and the comparison of MPC3 with various solvers is presented in Figs. 3(a) and 3(b). Another benefit of this MPC3 formulation is that previous solutions can be used to hot start the current iteration. In particular, if the time between control inputs is relatively small and the trajectory remains relatively similar on this scale (as is expected with MPC), the previous $\bm{\chi}$ can be used as an initial guess for the current coefficients and slack variable, significantly reducing the QP runtime.
@@ III-B Variation with Control Horizon
Further, the state dimension $q=6$ has been kept unchanged and the control horizon has been varied from $p=5$ to $p=20$ with an increment of $5$. The results presented in Figs. 3(c) and 3(d) particularly underscore the benefits of MPC3. Since the ICC solution is global over the time horizon of interest, the runtime or memory requirements of MPC3 formulation do not increase with a larger time horizon. While a larger approximation order $n$ may be necessary to capture the greater dynamic range over this time horizon, a small increase in $n$ is sufficient for these changes, which is shown to only marginally increase runtime.
@@ III-C Performance Comparison on an Edge Computer
The similar analysis is also carried out for Teensy 4.1111https://www.pjrc.com/store/teensy41.html development board to qualify the computational benefits of the proposed framework on an embedded hardware. The MATLAB® scripts used for the earlier analysis have been built and deployed on the Teensy board using automatic code generation. The results of this exercise are presented in Figure 4, follows a similar trend of run time and memory usage for various state dimensions and control horizon. Most importantly, as the control horizon increases, the RAM usage crosses the available on-board RAM of $512$ kb and the conventional MPC could no longer be implemented on the hardware. However, as mentioned earlier, the resource utilization does not increase with the control horizon within the MPC3 framework. This enables autonomous applications which were deemed infeasible due to high computational cost and memory usage.
@@ IV Application To Space Proximity Operations
The Transforming Proximity Operations and Docking System (TPODS) is a conceptual 1U CubeSat module, developed by the Land, Air and Space Robotics (LASR) laboratory at Texas A&M University [31, 32, 33]. The proposed MPC3 framework is utilized to enable robust and safe proximity operations of the TPODS module. The objective is to dock a chaser TPODS to a stationary target using a multi-state guidance algorithm. The TPODS module is equipped with Ultra-Wide-Band (UWB) radar range and monocular-vision camera sensors [33]. A unified pose estimator can be leveraged to fuse the range from stationary anchors with monocular vision data to produce consistent pose estimates [34]. The guidance logic for the docking scenarios considered in this paper is shown in Fig. 1.
@@ IV-A System Dynamics
Figure 5: Arrangement of UWB transceiver on TPODS module and stationary anchors. Since the anchors and UWB sensor are not mounted at respective centers of mass, the rotational and translation motion of the UWB sensor relative to the stationary anchors are coupled.
From the arrangement shown in Fig. 5, the motion governing equations for the chaser TPODS are [33]
$\bm{\ddot{\rho}_{ij}}$ $=\bm{\ddot{\rho}}+\bm{\dot{\omega}}\times\bm{P}_{1}^{i}+\bm{\omega}\times\left(\bm{\omega}\times\bm{P}_{1}^{i}\right)$ (43)
$I_{0}\bm{\dot{\omega}}$ $=I_{0}D(I_{1})^{-1}\left[\bm{N_{1}}-D^{\top}\bm{\omega}\times I_{1}D^{\top}\bm{\omega}\right]$ (44)
Here, $D$ is the rotation matrix that transforms vectors from TPODS reference frame $\bm{\hat{d}}$ to the mother-ship frame $\bm{\hat{c}}$, $\bm{\omega}$ is the rotational angular velocity of TPODS relative to the mother-ship, $I_{0}$ and $I_{1}$ are respective inertia tensors for the mother-ship and TPODS and $\bm{N_{1}}$ is the external torque applied to the TPODS. The CubeSat module has ten nozzles arranged in an ‘X’ configuration on each opposite face to enable 6-DOF [33].
For the simulation studies presented in this section, the linearized dynamics of the double-integrator in three translation axes are considered for the constrained optimization problem. Once the set of optimal control inputs is obtained for the selected control horizon, the control inputs at the first instance are determined and applied to the non-linear dynamics given by Eqs. (43) and (44), and the true states are propagated. The current states are fed to the multi-state guidance algorithm after an appropriate process noise has been injected into the velocity states. Finally, the desired values of the attitude and position computed by the guidance algorithm are fed to the optimization routine and the process continues.
Figure 6: Comparison of KOS constraint using discontinuous sate, repulsive potential and differential polytopic hull.
@@ IV-B Keep-Out-Constraint
One of the prominent safety requirements for the majority of real-world operations is Keep-Out-Constraints (KOC). Due to inherent uncertainties associated with localizing objects, it is necessary to account for the margin of safety. One of the ways this can be achieved is by prohibiting the presence of a chaser within a certain spherical or ellipsoidal volume. A few approaches to incorporate KOC in the MPC framework have been considered: a discontinuous augmented state [35], a repulsive pseudo-potential and differentiable collision detection of polytopic convex hulls [36]. While using the augmented state, the violation has to occur once for KOC to be effective. This limitation results in a steep correction and a large force demand and is not desirable as it causes intense structural stress and increased fuel consumption.
An alternative to the discontinuous augmented state is smooth pseudo potential which repulses the chaser when near to the target. This repulsive pseudo potential can be in the form $ar^{-l}$, where $r$ is the distance between the target and chaser. The tuning parameters $a,l>0$ depict the intensity of the repulsive action. These parameters, along with the cost weight factor, can be adjusted to get the desired radius of the KOC. One such example trajectory for $a=l=1$ is shown in Fig. 6. It is evident that the overall trajectory is considerably smoother and does not violate the KOC at all. It also achieves substantial fuel savings compared to the discontinuous augmented state.
Figure 7: Collision avoidance algorithm using differential collision detection of polytopic primitives
Although the KOC with repulsive potential works well for symmetric bodies, it does not result in an efficient utilization of the feasible space for axisymmetric elongated bodies. For such scenarios, differential collision detection for convex polytopes (DCOL) can provide a robust and efficient avenue to enforce KOC [36]. Collision detection via DCOL works by inflating polytopic convex hulls of the target and chaser by a scaling factor $s$. When $s$ is found to be greater than one, both bodies are separated. For a given set of polytopes, the minimum value of $s$ can be computed by solving a linear program with inequality constraints stemming from the fact that the intersection point is a member of both inflated polytopes.
Once the collision is detected, corrective action needs to be taken such that the chaser moves away from the target. For the analysis presented in this paper, the target is considered stationary and the attitude of the chaser is not varied during collision avoidance maneuvers. Hence, the general direction in which the separation is achieved can be inferred from the relative change of the $s$ with respect to the position of the chaser. For the MPC framework, collision avoidance can be achieved with the process shown in Fig. 7 [37]. The optimization problem given by Eq. (34) is first solved without any additional collision avoidance constraints. Next, the value of $s$ for all instances within the control horizon is computed.
If the value of $s$ for any instance within the control horizon is less than a predefined soft limit $s_{\text{thr}}$, the partial derivative $\frac{\partial s}{\partial\bm{r}_{c}}$ is computed and the optimization problem is modified as
$\min_{\bm{u},\mathop{}\!\mathrm{d}{\bm{u}},\bm{\varepsilon}}\qquad$ $\lvert|\bm{u}+\mathop{}\!\mathrm{d}{\bm{u}}|\rvert_{2}+\lvert|\bm{y}-\bm{y}_{\text{tar}}|\rvert_{2}+\rho\varepsilon^{2}$ (45)
@@
$\bm{u}_{\max}$ $\geq\lvert\bm{u}+\mathop{}\!\mathrm{d}{\bm{u}}\rvert$ (47)
$s_{\text{thr}}$ $\leq s+\frac{\partial s}{\partial\bm{r}_{c}}\frac{\partial\bm{r}_{c}}{\partial\left(\bm{u}+\mathop{}\!\mathrm{d}{\bm{u}}\right)}+\rho\varepsilon$ (48)
The modified input sequence which ensures a collision-free trajectory is then applied to the chaser and the process is repeated until the chaser is within a specified radius of the target. As shown in Fig. 6, collision avoidance with DCOL results in a smooth trajectory with the least deviation from the intended optimal trajectory among all approaches analyzed. In addition to the least energy consumption, the collision avoidance approach based on the polytopic hulls of respective bodies prevents the optimization problem from becoming infeasible due to geometrical complexities [36].
@@ IV-C Optimal Trajectories for Docking
To validate the proposed framework, the orientation of the docking face on the target and the initial pose of the chaser is selected such that the extremal trajectory of the chaser passes through the target. Individual translation velocities are restricted to $0.02$ m/s with a softness factor of $0.1$ and the control forces are restricted to $10$ mN in each axis with a softness factor of $0.01$. The soft target for scaling factor $s_{\text{thr}}$ is picked as $1.5$ to ensure robustness against pose estimation errors. The collision detection and avoidance is switched on when the chaser is within $0.4$ m radius of the target and switched off during the final phase of docking. The attitude controller acts independently and leverages errors in desired and current quaternion to drive the chaser to the desired orientation during various guidance modes [34].

### v2

@@ II-C State and Input Constraints
With these modifications, the QP of Eq. (12) is written with equality and inequality constraints as
$\begin{split}\min_{\bm{\chi}}J&=\frac{1}{2}\bm{\chi}^{\top}H\bm{\chi}+\bm{f}^{\top}\bm{\chi}\\ \text{s.t.}\quad A_{\text{eq}}\bm{\chi}&=\mathbf{b}_{\text{eq}},\quad A\bm{\chi}\leq\mathbf{b}\end{split}$ (14)
where, with some abuse of notation,
$\bm{\chi}$ $=\begin{bmatrix}\bm{\alpha}\\ \varepsilon\end{bmatrix}\in\mathbb{R}^{(n+2)\times 1}$ (15a)
$H$ $=\begin{bmatrix}H&\bm{0}_{n+1,1}\\ \bm{0}_{1,n+1}&\rho\end{bmatrix}\in\mathbb{R}^{(n+2)\times(n+2)}$ (15b)
$\bm{f}$ $=\begin{bmatrix}\bm{f}\\ 0\end{bmatrix}\in\mathbb{R}^{(n+2)\times 1}$ (15c)
$A_{\text{eq}}$ $=\begin{bmatrix}\bm{\gamma}(-1)&0\\ \bm{\beta}(-1)&0\end{bmatrix}\in\mathbb{R}^{2\times(n+2)}$ (15d)
$\mathbf{b}_{\text{eq}}$ $=\begin{bmatrix}0\\ 0\end{bmatrix}\in\mathbb{R}^{2\times 1}$ (15e)
$A$ $=\begin{bmatrix}\mathcal{T}(\bm{\tau})&-V_{u}\\ -\mathcal{T}(\bm{\tau})&-V_{u}\\ \bm{\beta}(\bm{\tau})&-V_{x^{\prime}}\\ -\bm{\beta}(\bm{\tau})&-V_{x^{\prime}}\end{bmatrix}\in\mathbb{R}^{4(n+1)\times(n+2)}$ (15f)
$\mathbf{b}$ $=\begin{bmatrix}u_{\text{max}}^{\prime}\\ -u_{\text{min}}^{\prime}\\ x_{\text{max}}^{\prime}-x^{\prime}(-1)\\ -x_{\text{min}}^{\prime}+x^{\prime}(-1)\end{bmatrix}\in\mathbb{R}^{4(n+1)\times 1}$ (15g)
Here, $\rho$ is the weight factor associated with the slack variable $\varepsilon$, $u_{\text{max}}^{\prime}/u_{\text{min}}^{\prime}$ and $x_{\text{max}}^{\prime}/x_{\text{min}}^{\prime}$ are the constraints on the control and velocity, and $V_{u}$ and $V_{x^{\prime}}$ are used to regulate the softness of the constraints. A comprehensive comparison of MPC3 with discrete MPC along with dimensions of the optimization program is given in Table I.
@@ III Performance Comparison
The constrained optimization problem, now formulated in the computational domain as a QP, is solved with MATLAB®’s quadprog using the active-set algorithm and warm-start enabled. Importantly, note that of the matrices and vectors defined in the previous section, $H$, $A_{\text{eq}}$, $\mathbf{b}_{\text{eq}}$, and $A$ remain unchanged, regardless of time horizon or initial conditions. Only $\bm{f}$ and $\mathbf{b}$ need to be updated at each step, resulting in an efficient formulation. The coefficients and slack variable are found by minimizing the objective function of Eq. (14), and the instantaneous control is calculated as $u(-1)=\bm{T}(-1)\bm{\alpha}$. ICC, or any other integrator, can then be used to propagate the true (generally nonlinear) dynamics until reaching the next control input time.
@@ III-A Variation with State Dimension
To establish the computational gains of the proposed formulation, extensive simulation studies have been conducted and their results are presented in Fig. 2. For the first analysis, the control horizon has been kept fixed at $p=5$ samples and the dimension of the state $q$ has been varied. The average run time and memory requirement of a single optimization routine is recorded and the comparison of MPC3 with various solvers is presented in Figs. 2(a) and 2(b). Another benefit of this MPC3 formulation is that previous solutions can be used to hot start the current iteration. In particular, if the time between control inputs is relatively small and the trajectory remains relatively similar on this scale (as is expected with MPC), the previous free variables $\bm{\chi}$ can be used as an initial guess for the current coefficients and slack variable, significantly reducing the time to solve QP.
@@ III-B Variation with Control Horizon
Further, the state dimension $q=6$ has been kept unchanged and the control horizon has been varied from $p=5$ to $p=20$ with an increment of $5$. The results presented in Figs. 2(c) and 2(d) particularly underscore the benefits of MPC3. Since the ICC solution is global over the time horizon of interest, the runtime or memory requirements of MPC3 formulation do not increase with a larger time horizon. While a larger approximation order $n$ may be necessary to capture the greater dynamic range over this time horizon, a small increase in $n$ is sufficient for these changes, which is shown to only marginally increase runtime.
@@ III-C Performance Comparison on an Edge Computer
A similar analysis is also carried out for Teensy 4.1111https://www.pjrc.com/store/teensy41.html development board to qualify the computational benefits of the proposed framework on an embedded hardware. The MATLAB® scripts used for the earlier analysis have been built and deployed on the Teensy board using automatic code generation. The results of this exercise are presented in Figure 3, follows a similar trend of run time and memory usage for various state dimensions and control horizon. Most importantly, as the control horizon increases, the RAM usage crosses the available on-board RAM of $512$ kb and the conventional MPC could no longer be implemented on the hardware. However, as mentioned earlier, the resource utilization does not increase with the control horizon within the MPC3 framework. This enables autonomous applications which were deemed infeasible due to high computational cost and memory usage.
@@ IV Application To Space Proximity Operations
The Transforming Proximity Operations and Docking System (TPODS) is a conceptual 1U CubeSat module, developed by the Land, Air and Space Robotics (LASR) laboratory at Texas A&M University [24, 25]. The proposed MPC3 framework is utilized to enable robust and safe proximity operations of the TPODS module. The objective is to dock a chaser TPODS to a stationary target using a multi-state guidance algorithm. The TPODS module is equipped with Ultra-Wide-Band (UWB) radar range and monocular-vision camera sensors [25]. A unified pose estimator can be leveraged to fuse the range from stationary anchors with monocular vision data to produce consistent pose estimates. The guidance logic for the docking scenarios considered in this paper is shown in Fig. 4.
@@ IV-A System Dynamics
The UWB sensor is not mounted on the respective centers of mass for the TPODS module, resulting in a coupled rotational and translational motion of the UWB sensor relative to the stationary anchors[26]. For the simulation studies presented in this section, the TPODS dynamics are linearized into a double integrator in three translation axes. Once the set of optimal control inputs is obtained for the selected control horizon, the control inputs at the first instance are determined and applied to the non-linear dynamics, and the true states are propagated. The current states are fed to the multi-state guidance algorithm after an appropriate process noise has been injected into the velocity states. Finally, the desired values of the attitude and position computed by the guidance algorithm are fed to the optimization routine and the process continues.
@@ IV-B Keep-Out-Constraint
One of the prominent safety requirements for the majority of real-world operations is Keep-Out-Constraints (KOC). Due to inherent uncertainties associated with localizing objects, it is necessary to account for the margin of safety. A differential collision detection for convex polytopes (DCOL) can provide a robust and efficient avenue to enforce KOC [27]. Collision detection via DCOL works by inflating polytopic convex hulls of the target and chaser by a scaling factor $s$[26]. When $s$ is found to be greater than one, both bodies are separated. For a given set of polytopes, the minimum value of $s$ can be computed by solving a linear program with inequality constraints stemming from the fact that the intersection point is a member of both inflated polytopes.
Once the collision is detected, corrective action needs to be taken such that the chaser moves away from the target. For the analysis presented in this paper, the target is considered stationary and the attitude of the chaser is not varied during collision avoidance maneuvers. Hence, the general direction in which the separation is achieved can be inferred from the relative change of the $s$ with respect to the position of the chaser. The optimization problem given by Eq. (14) is first solved without any additional collision avoidance constraints. Next, the value of $s$ for all instances within the control horizon is computed.
@@ IV-C Optimal Trajectories for Docking
To validate the proposed framework, the orientation of the docking face on the target and the initial pose of the chaser is selected such that the extremal trajectory of the chaser passes through the target. Individual translation velocities are restricted to $0.02$ m/s with a softness factor of $0.1$ and the control forces are restricted to $10$ mN in each axis with a softness factor of $0.01$. The soft target for scaling factor $s_{\text{thr}}$ is picked as $1.5$ to ensure robustness against pose estimation errors. The collision detection and avoidance is switched on when the chaser is within $0.4$ m radius of the target and switched off during the final phase of docking. The attitude controller acts independently and leverages errors in desired and current quaternion to drive the chaser to the desired orientation during various guidance modes.